OpenStudy (exoexoexo):

three numbers whose common difference is 5 are in an arithmetic sequence. if the first number is left unchanged, and 1 is subtracted from the second, and 2 is added to the third, the resulting three numbers are in a geometric sequence. find the three original numbers and the geometric sequence.

OpenStudy (retireed):

Did you get an answer yet? I think I solved it, if you want help solving it.

OpenStudy (exoexoexo):

I still can't figure it out

OpenStudy (sshayer):

let the numbers be a-5,a,a+5 new numbers are a-5,a-1,a+5+2 or a-5,a-1,a+7 because they are in G.P \[\left( a-1 \right)^2=\left( a-5 \right)\left( a+7 \right)\] calculate a and then the numbers

OpenStudy (exoexoexo):

@sshayer I don't understand why you have to square (a-1) can you explain?

OpenStudy (sshayer):

if a,b,c are in G.P then \[\frac{ b }{ a }=\frac{ c }{ b }~or~b^2=ac\]

OpenStudy (exoexoexo):

I got \[a ^{2}-2a+1 = a ^{2}+2a+35\]

OpenStudy (sshayer):

good take all the terms with a and a^2 on one side and constant terms on other side and find a

OpenStudy (sshayer):

it is -35

OpenStudy (sshayer):

-5*7=-35

OpenStudy (exoexoexo):

oops

OpenStudy (exoexoexo):

9=a

OpenStudy (sshayer):

now find the original numbers

OpenStudy (exoexoexo):

4,9,14

OpenStudy (sshayer):

correct

OpenStudy (sshayer):

now find the numbers of G.P

OpenStudy (exoexoexo):

4,8,16

OpenStudy (sshayer):

correct well done.

OpenStudy (exoexoexo):

thank you

OpenStudy (sshayer):

yw